Find the domain of the function f(x)= 1+cos^2x/1-sin^2x

Grade:11

1 Answers

MuraliKrishna Medavaram
10 years ago
f(x) = 1+cos^2x/1-sin^2x = 1+cos^2x/cos^2x
=sec^2x+1
and hence domain of f(x) is domain of sec^2x
which is

Thanks and Regards,
M.MURALIKRISHNA
askIITians faculty